OpenStudy (anonymous):

IDK how to figure it out..but i got a clue: The line which you folded is called the line of symmetry. This divides the parabola into two matching halves or mirror images. The equation of the line is x = ___ because it is vertical. Look to see where your folded line crosses the x-axis. Choose the equation below. A. x = -2 B. x = 2 C. x = 0 D. None of the above

OpenStudy (anonymous):

I think it's C

OpenStudy (ash2326):

What's the equation of the parabola?

OpenStudy (anonymous):

would be nice to have either the equation or a picture

OpenStudy (anonymous):

id even have a equation or picture..this is all tehy gave me.

OpenStudy (anonymous):

Look to see where your folded line crosses the x-axis.

OpenStudy (ash2326):

If your curve is y=x^2 , then line of symmetry is x=0

OpenStudy (anonymous):

you must be given something right? there is no information here about what the parabola is

OpenStudy (anonymous):

x=-b/2a is the line of symmetry

OpenStudy (anonymous):

look all i got was this now in previous questions i got some equations and idk if they ar erefering to tht but they didn't give me one in this question @satellite73

OpenStudy (anonymous):

they have to be referring to this question from the last equation

OpenStudy (anonymous):

it must be about the previous questions, because otherwise this makes no sense the axis of symmetry can be any vertical line look at the previous one and let us know what it is

OpenStudy (anonymous):

the last equation was y = x^2+4x+1

OpenStudy (anonymous):

ahhh now we are getting somewhere \[-\frac{b}{2a}=-\frac{4}{2}=-2\]

OpenStudy (anonymous):

so answer is \[x=-2\] as you suspected

OpenStudy (anonymous):

oh okay yea i wuld have got tht question rong....thnks u guys for ur hlp...u all get medals

OpenStudy (anonymous):

yw